# Federated Anonymous Blocklisting across Service Providers and its Application to Group Messaging

### v1

@@ 6 Implementation
We have implemented the proposed FAB scheme in $\sim 3000$ lines of Rust code. [^1] Available at https://github.com/SDABIS/mls_fab Table 1 shows the concrete algorithms employed for each of the cryptographic primitives our construction depends on. All zk-SNARK operations are handled by the arkworks ecosystem [arkworks].
Table 1: Algorithms employed for each operation of the protocol.
@@
Hash Function Poseidon [poseidon]
@@
zk-SNARK scheme Groth16 [groth16]
In our implementation we employ a Merkle Tree as negative accumulator to represent the list of banned users. This structure allows for logarithmic complexity in membership verification, which is particularly relevant as verification is executed inside the Block circuit [zcash].
Merkle Trees only allow for membership proofs, whereas we only require non-membership proofs. To that end, we construct a Complementary Merkle Tree [promises]: starting with the set of all possible pseudonyms, it is divided into intervals $[a,b)$ with a breakpoint at all of the banned pseudonyms. Every leaf in the Complementary Merkle Tree represents one of said intervals and contains $H(a||b)$. To prove non-membership of $ps$, users must prove that there exists a leaf at position $i$ with content $H(a||b)$ such that $a\leq ps<b$, that is, $\overline{w_{x}}=(a,b,i,ps)$.
@@ 6.1 Integration into MLS
We also provide a concrete application of the theoretical concepts of FAB to demonstrate its real-world usefulness. In particular, we apply FAB to the context of instant messaging groups. Each realm represents a different messaging group with its own blocklist. We employ the Messaging Layer Security [mls] framework for our proof-of-concept implementation.
Every FAB operation is expressed in the terms of MLS. The functions that modify the state of the realm $\mathsf{Block}$, $\mathsf{Unblock}$ and establishing trust relations— are performed by the proposal-and-commit paradigm of MLS: users can propose this modifications to the group, which are later committed.
@@
We note that establishing consensus among group members about which pseudonyms should be added is outside the scope of this work. There exist other works that are concerned with authorisation in MLS groups that are compatible with our environment [cgka_fa, a_cgka].
@@ 6.2 Evaluation
In this Section, we evaluate the performance of our proposed FAB scheme. For comparison, benchmarks for the state-of-the-art SNARKBlock and ALPACA schemes are also included. SNARKBlock is configured with chunk sizes of 1024 pseudonyms and a buffer of 14 chunks of size 16. The measurements for all three schemes were obtained in the same execution environment, using the code provided in their respective repositories [snarkblock-impl, alpaca-impl].
@@
Figure 4: Performance evaluation of our FAB scheme and comparison with state-of-the-art Anonymoust Blocklisting scheme SNARKBlock and ALPACA. We represent our FAB scheme for multiple amounts of trusted realms. The measurements combine the cost of $R^{auth}$ and $R^{ban}$, the latter being executed once for every trusted realm.
@@ 7 Discussion
In this Section, we comment on real-world aspects of our construction.
@@ 7.1 Comparison with other Anonymous Blocklisting schemes
Although related, our FAB scheme is based on significantly different assumptions about the structure of service providers than other AB works such as [blac, snarkblock, alpaca]. These works assume a single service provider with high traffic; indeed, networks with thousands of daily blocks such as Wikipedia or Reddit are often used as examples.
In contrast, our model considers a high number of Realms with less traffic and thus more sparse blocks. Throughout this document we have used messaging groups to illustrate our envisioned scenario. However, FAB schemes are also applicable to big Service Providers that are divided into sub-communities, such as Reddit or Discord. Instead of using a single blocklist for the service, each of the communities could represent a different realm in a FAB scheme.
@@
Table 2 compares the algorithmic complexity of our scheme with the AB schemes SNARKBlock and ALPACA. All of our measurements scale linearly with the number of realms, with Auth times also being affected by the depth of the blocklist Merkle Tree. This is a direct improvement over SNARKBlock, whose performance increases linearly with the blocklist size. Conversely, the complexity of ALPACA is constant: as shown in Figure 4(a), its authentication cost is roughly equivalent to a FAB authentication with 20 trusted realms.
@@
7.2 Limitations and alternative constructions
We now discuss the limitations of our scheme and propose alternative constructions to address them.
Firstly, our scheme does not allow proof reusability. In our construction, proofs for the $R^{ban}$ relation also include the pseudonym of the target realm. This ensures that the $\pi^{ban}_{i}$ proofs were created with the same credential than their corresponding $\pi^{auth}$ proof. Unfortunately, that means that $\pi^{ban}$ cannot be reused for authentication in other target realms. Alternatively, aggregable zk-SNARKs could be employed [zkcreds] to ensure that multiple proofs share a common input. This introduces the additional cost of aggregating the proofs, but may represent an acceptable trade-off if the $\pi^{ban}$ proofs are recomputed often.
We also address the pseudonym lifetime. Whereas in AB schemes each pseudonym corresponds only to a single message, ours is linked to authentication and thus is reused for every interaction inside a given realm. This may be an issue depending on the security requirements of the application. As a solution, user credentials could be expanded to have $n$ identities instead of only one and thus $n$ possible pseudonyms per realm. Proving the $R^{block}$ relation would involve showing that none of the user’s pseudonyms has been banned, which would increase the cost by a factor of $n$. Alternatively, batch proofs [reckle] could be used to limit the cost increase.

### v2

@@ 6 Implementation
We have implemented the proposed FAB scheme in $\sim 3000$ lines of Rust code. [^1] Available at https://github.com/SDABIS/mls_fab Table 1 shows the concrete algorithms employed for each of the cryptographic primitives our construction depends on. All zk-SNARK operations are handled by the arkworks ecosystem [3].
Table 1: Algorithms employed for the implementation.
@@
Hash Function Poseidon [10]
@@
zk-SNARK scheme Groth16 [11]
In our implementation we employ a Merkle Tree as negative accumulator to represent the list of banned users. This structure allows for logarithmic complexity in membership verification, which is particularly relevant as verification is executed inside the Block circuit [16]. Merkle Trees only allow for membership proofs, whereas we only require non-membership proofs. To that end, we construct a Complementary Merkle Tree [30]: starting with the set of all possible pseudonyms, it is divided into intervals $[a,b)$ with a breakpoint at all of the banned pseudonyms. Every leaf in the Complementary Merkle Tree represents one of said intervals and contains $H(a||b)$. To prove non-membership of $ps$, users must prove that there exists a leaf at position $i$ with content $H(a||b)$ such that $a\leq ps<b$, that is, $\overline{w_{x}}=(a,b,i,ps)$.
We additionally include in the implementation a Realm Directory as described in Figure 2, instantiated as a REST API web server. Realms upload their states to the Directory, which are downloaded by clients for the authentication.
@@ 6.1 Integration into MLS
We also provide a concrete application of the theoretical concepts of FAB to demonstrate its real-world usefulness. In particular, we apply FAB to the context of instant messaging groups. Each realm represents a different messaging group with its own blocklist. We employ the Messaging Layer Security [6] framework for our implementation.
Every FAB operation is expressed in the terms of MLS. The functions that modify the state of the realm $\mathsf{Block}$, $\mathsf{Unblock}$ and establishing trust relations— are performed by the proposal-and-commit paradigm: users can propose this modifications to the group, which are later committed.
@@
We note that establishing consensus among group members about which pseudonyms should be added is outside the scope of this work. There exist other works that are concerned with authorisation in MLS groups that are compatible with our environment [19, 4].
@@ 6.2 Evaluation
In this Section, we evaluate the performance of our proposed FAB scheme. For comparison, benchmarks for the state-of-the-art SNARKBlock [28] and ALPACA [20] schemes are also included. SNARKBlock is configured with chunk sizes of 1024 pseudonyms and a buffer of 14 chunks of size 16. The measurements for all three schemes were obtained in the same execution environment, using the code provided in their respective repositories [29, 18].
@@
Figure 12: Performance evaluation of our FAB scheme and comparison with state-of-the-art Anonymous Blocklisting scheme SNARKBlock and ALPACA. We represent our FAB scheme for multiple amounts of trusted realms. The measurements combine the cost of $R^{auth}$ and $R^{block}$, the latter being executed once for every trusted realm.
@@ 7 Discussion
We now comment on real-world aspects of our construction.
@@ 7.1 Comparison with other Anonymous Blocklisting schemes
Table 2 compares the algorithmic complexity of our scheme with the AB schemes SNARKBlock and ALPACA. All of our measurements scale linearly with the number of realms, with Auth times being affected by the depth of the blocklist Merkle Tree. This is an improvement over SNARKBlock, whose performance increases linearly with the blocklist size. Conversely, the complexity of ALPACA is constant: as shown in Figure 12(a), its authentication cost is roughly equivalent to a FAB authentication with 20 trusted realms.
@@
7.2 Advantages and limitations
Although related, our FAB scheme is based on significantly different assumptions about the structure of service providers than other AB works such as [35, 28, 20]. These works assume a single service provider with high traffic; indeed, networks with thousands of daily blocks such as Wikipedia or Reddit are often used as examples.
In contrast, our model considers a high number of comparatively smaller realms. Throughout this document we have used messaging groups to illustrate our envisioned scenario. However, the FAB scheme is also applicable to big Service Providers that are divided into sub-communities, such as Reddit or Discord. Instead of using a single blocklist for the service, each of the communities could represent a different realm in a FAB scheme. This is advantageous as our FAB scheme outperforms other AB schemes in scenarios in which blocks are issued frequently, since it does not require synchronisation. However, the FAB scheme introduces a new linear scaling: the number of trusted realms. If this number is too high, then FAB may become less efficient than traditional AB schemes. Figure 12(a) analyses up to 100 trusted realms, resulting in approximately 15 seconds for authentication.
The trust requirements also different for both schemes. In AB, users have no power to insert or remove pseudonyms from the blocklist, and rely on the Service Provider administrators for these activities. In contrast, FAB schemes allow realm creators to dynamically decide whether or not they trust a particular realm, allowing for a more fine-grained definition of trust. This choice on whether or not to trust on other realms could be based on their similarity of rules of conduct, level of moderation or track record. Admittedly, evaluating other realms to decide if they are trustworthy requires an additional effort in group administration. We leave for future work whether or not it is possible to merge or aggregate different realms, which would minimise the cost of such administration tasks.
Lastly, the scheme presented in this work offers Unlinkability across different realms, but not inside the same realm. Indeed, a user must employ the same pseudonym for multiple authentication attempts inside a realm. Traditional AB schemes would be more appropiate in scenarios in which every individual message in a given realm needs to be unlinkable. Our scheme could be modified to provide limited Unlinkability inside a given realm by expanding user credentials to have $n$ identities $x_{0},...x_{n}$ instead of only one $x$. The modification would allow users to have $n$ possible pseudonyms in a given realm. Unfortunately, this would significantly affect the performance of the $R^{block}$ relation, as it would involve showing that none of the user’s $n$ pseudonyms has been blocked, increasing the cost by a factor of $n$. The cost increase could be mitigated by the inclusion of batch proofs [26], which would allow users to efficiently prove non-membership of their $n$ identities.
@@ Acknowledgements
This work has been funded by the European Regional Development Fund (ERDF) through the EU Interreg VI-A Spain-Portugal (POCTEP) 2021-2027 Programme, project ”Quantum IBER_IA: Impulso estratégico de las capacidades en tecnologías cuánticas e inteligencia artificial en el espacio ibérico transfronterizo”. D.S. acknowledges support from Xunta de Galicia and the European Union (European Social Fund - ESF) scholarship [ED481A-2023-219]. We also acknowledge support from the Xunta de Galicia and the European Union (FEDER Galicia 2021-2027 Program) Ref. ED431B 2024/21, CITIC ED431G 2023/0.